OpenStudy (loser66):

\(in~~Z_{24},~~find~~x~~if~~x^n=0\) please, help

OpenStudy (loser66):

@ganeshie8

ganeshie8 (ganeshie8):

wouldn't x=0 work ?

OpenStudy (loser66):

non trivial, please

OpenStudy (anonymous):

I sense a disturbance in the force...

ganeshie8 (ganeshie8):

lol dumb q : what do we know about \(n\) ?

OpenStudy (loser66):

n in Z+

OpenStudy (anonymous):

What is \(\Large \mathbb{Z}_{24}\) taken to be? A ring?

OpenStudy (loser66):

yes, @Kurt

OpenStudy (loser66):

I have problem with the net... OMG.... I need know how to work out.... Please.. leave the guidance !!

OpenStudy (anonymous):

...and, what? Are we looking for an element of \(\large \mathbb{Z}_{24}\) that is always zero when you raise it to any positive integer?

OpenStudy (loser66):

I think so

OpenStudy (loser66):

not any, some n

OpenStudy (anonymous):

for some n? Okay, that simplifies things a bit.

OpenStudy (anonymous):

Hmm.. @ganeshie8 ? :D

ganeshie8 (ganeshie8):

does this work ? \[x=2^a3^b\] where \(a\ge 1\) and \(b\ge 1\)

OpenStudy (anonymous):

yes it does, @ganeshie8 now... let's prove it! :D

OpenStudy (loser66):

yes, I think so, but what is the logic? I mean how to get that?

OpenStudy (anonymous):

time for a little Kurt-magic? :) haha, kidding :>

OpenStudy (loser66):

I have school now, but I do need know how to get it. Please. I appologize for not being here to get help.

OpenStudy (anonymous):

btw, @ganeshie8 \[\Large x = 2^a \cdot 3^b \cdot \color{blue}k\] seems to work, for all \[\Large a,b,k \in \mathbb{Z}^+\]

ganeshie8 (ganeshie8):

that looks complete now :)

OpenStudy (anonymous):

In English, L66, x must be divisible by both 2 and 3. The plan has two phases, first, prove that if x is divisible by 2 and 3, then eventually, multiplying it by itself will yield a multiple of 24. Second, show that IF x is indivisible by either 2 or 3, then it will never yield a multiple of 24 no matter how many times you multiply it by itself.

OpenStudy (anonymous):

Okay, phase 1. Suppose x is divisible by both 2 and 3. That means that x is divisible by 6. So... \(\large x = 6k\) Then take n = 3. \[\Large x^3 = 6^3k^3 = 216k^3 = 24\cdot 9\cdot k^3 \equiv 0 (\text{mod} \ 24)\]

OpenStudy (anonymous):

Phase 2a: Suppose x is odd. That is to say, it is indivisible by 2. It's simple enough to show that any integer power of an odd number will be odd. (Try to do that yourself: use induction!) So, if x is odd, then x^n will also be odd, and therefore, we cannot have \[\Large x^n = 24k = 2\cdot 12k \]

OpenStudy (anonymous):

x^n will never be divisible by 24, no matter what positive integer value n takes.

OpenStudy (anonymous):

Phase 2b requires you to show that if x is indivisible by 3, then it also cannot be raised to any power such that it will be divisible by 24. Use a similar approach to phase 2a. Good luck!

OpenStudy (anonymous):

And one non-essential hint that may well make your life easier: All integers that are not divisible by 3 may be written as \[\Large 3k \pm 1\] Where k is an integer. Best of luck :D

ganeshie8 (ganeshie8):

nice

OpenStudy (loser66):

??? if \(x^n\equiv 0\) (mod 24) , then \(24| x^n\) 24 =2*2*2*3 \(x^n =x*x*......*x~(ntimes)\) hence 2*2*2*3|x*x*......*x which leads to 2|x and 3|x or 6|x the element x in \(Z_{24} \) satisfy x =6k are {0,6,12,18} Does it work?